Which is a solution for the following system of equations?
y = x^2
y=-4x + 12

Answer by stanbon(75887)

Which is a solution for the following system of equations?
y = x^2
y=-4x + 12
----------------------
Substitute and solve for "x":
x^2= -4x+12
---
x^2+4x-12 = 0
x^2+6x-2x-12 = 0
x(x+6)-2(x+6) = 0
(x+6)(x-2) = 0
x = -6 or x = 2
------------------------------------
If x = -6, y = 36
-----
If x = 2, y = 4
======================
Cheers,
Stan H.